OpenStudy (anonymous):

convert following logarithms to exponential form ln(4x)=9.6 log5 1/25=-2 5 is lower than word log

OpenStudy (anonymous):

e^9.6 = 4x

OpenStudy (anonymous):

5^(-2) = 1/25

OpenStudy (anonymous):

Well for the first one just raise everything to the e. 4x=e^9.6 Then for the second one 1/25=5^-2

OpenStudy (anonymous):

how about log8 64=2 the 8 is lower

OpenStudy (anonymous):

8^2 = 64

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome